OpenStudy (anonymous):

Multiply (6x + 7) (3x +2)

OpenStudy (anonymous):

FOIL-First Outer Inner Last Multiply First terms 6x\[\times\]3x=18x^2 Multiply Outer terms 6x\[\times\]2=12x Multiply Inner terms 7\[\times\]3x=21x Multiply Last terms 7\[\times\]2=14 Combine18x^2+12x+21x+14=18x^2+33x+14

OpenStudy (anonymous):

ty